• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 13:32:41 on April 1,2008

    Entry number 229343

    keyword=target status: loop 2 temp sensor problems

    Target is stable except there are problems with temperature sensors for loop2: Cell_aft was bad from beginning (yesterday), Htr_aft and Htr_fore started to fluctuate today. Htr_fore is used for PID control.
    Since it is fluctuation, I had to change the Heater to be in manual mode. Target group (Chris and Dave) are aware of the problem.


    FIGURE 1

    FIGURE 2